John Fleming, chief administrator for Valley View Hospital, is concerned about the costs for tests in the hospitals lab Charges for lab tests are consistently higher at Valley View than at other hospitals and have resulted in many complaints. Also, because of strict regulations on amounts reimbursed for lab tests, payments received from insurance companies and governmental units have not been high enough to cover lab costs 10 points Mr. Fleming has asked you to evaluate costs in the hospitals lab for the past month. The following information is available eBook a. Two types of tests are performed in the lab-blood tests and smears. During the past month, 1,800 blood tests and 2,400 smears were performed in the lab. b. Small glass plates are used in both types of tests. During the past month, the hospital purchased 12,000 plates at a Print cost of $56,400.1,500 of these plates were unused at the end of the month; no plates were on hand at the beginning of the month. c. During the past month, 1,150 hours of labor time were recorded in the lab at a cost of $21,850. References d. The labs variable overhead cost last month totaled $7,820 Valley View Hospital has never used standard costs. By searching industry literature, however, you have determined the following nationwide averages for hospital labs: Plates: Two plates are required per lab test. These plates cost $5.00 each and are disposed of after the test is completed. Labor: Each blood test should require 0.3 hours to complete, and each smear should require 0.15 hours to complete The average cost of this lab time is $20 per hour. Overhead: Overhead cost is based on direct labor-hours. The average rate for variable overhead is $6 per hour Required 1. Compute a materials price variance for the plates purchased last month and a materials quantity variance for the plates used last month 2. For labor cost in the lab: a. Compute a labor rate variance and a labor efficiency variance b. In most hospitals, one-half of the workers in the lab are senior technicians and one-half are assistants. In an effort to reduce costs, Valley View Hospital employs only one-fourth senior technicians and three-fourths assistants. Would you recommend that this policy be continued? 3-a. Compute the variable overhead rate and efficiency variances. 3-b. Is there any relation between the variable overhead efficiency variance and the labor efficiency variance? Complete this question by entering your answers in the tabs below. Req 1 Req 2A Req 2B Req 3A Req 3B Compute a materials price variance for the plates purchased last month and a materials quantity variance for the plates used last month. (Indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance). Input all amounts as positive values.) Show less Materials price variance Materials quantity variance Req Req 2A>

Please show step by step on how to do Req 1 , 2A , and 3A. Thank you.
Answer #1
1) Material price variance = Actual quantity*(Actual price-Standard price) = 12000*(56400/12000-5) = $          3,600 Favorable
Material quantity variance = Standard price*(Actual quantity-Standard quantity) = 5*(10500-4200*2) = $       10,500 Unfavorable
2) a) Labor rate variance = Actual labor hours*(Actual rate-Standard rate) = 1150*(21850/1150-20) = $          1,150 Favorable
Labor efficiency variance = Standard rate*(Actual hours-Standard hours) = 20*(1150-1800*0.3+2400*0.15) = $       19,400 Unfavorable
b) No, the policy should not be continuted. The reason is that the labor efficiency has gone down,
which, has resulted in an unfavorable variance
that has more than offset the favorable rate
variance.
3-a) VOH rate variance = Actual labor hours*(Actual VOG rate-Standard VOH rate) = 1150*(7820/1150-6) = $             920 Unfavorable
VOH efficiency variance = Standard VOH rate*(Actual hours-Standard hours) =6*(1150-1800*0.3+2400*0.15) = $          5,820 Unfavorable
3-b) Yes, both are based on labor hours--actual labor
hours and standard labor hours.